The greenhouse I am building for my wife is 6 feet longer than it is wide. Write an algebraic expression for the perimeter if the length is expressed as X. If X equals 20, what is the perimeter?
Found 2 solutions by praseenakos@yahoo.com, stanbon:
Answer by praseenakos@yahoo.com(507)

QUESTION:
The greenhouse I am building for my wife is 6 feet longer than it is wide. Write an algebraic expression for the perimeter if the length is expressed as X. If X equals 20, what is the perimeter?
ANSWER:

Given,

Length = x feet

Width = x -6 (because length is 6 ft longer than width means, width is 6 feet shorter than length.)


Perimeter of a rectangle is = 2 ( length + width)


==> perimeter = 2 ( x + x -6 )

= 2 ( 2x - 6 )

= 2* 2x - 2 * 6

= 4x - 12


If we denote perimeter by p, the algebraic expression for perimeter is,


p = 4x - 12


It is given that x = 20


==> p = 4 * 20 - 12

= 80 - 12

= 68


perimeter = 68 sqare feet.




Hope you understood.


Regards,
praseenakos@yahoo.co.in





Answer by stanbon(48564)

The greenhouse I am building for my wife is 6 feet longer than it is wide. Write an algebraic expression for the perimeter if the length is expressed as X. If X equals 20, what is the perimeter?
---------------------
length = x = 20
width = x-6= 14
Perimeter = 2*length + 2*width
Perimeter = 2*20 + 2*14=40+28=68 feet
Cheers,
Stan H.